# Sharing is CAIRing: Characterizing Principles and Assessing Properties of Universal Privacy Evaluation for Synthetic Tabular Data

### v1

@@ 5 Application of CAIR
The overall CAIR scores are given by the means with the standard error in Figure 2 where the best metric clearly is $\epsilon$-identifiability with a score of $3.56\pm 0.15$. The metric scores well in most dimensions and notably has some high scores in all four main categories. In contrast, the most closely related metric to $\epsilon$-identifiability, IDR, also performs well, but is especially penalized in the Applicability category for requiring information that is difficult to obtain such as equivalence classes of the true population. Consequently, IDR gets a considerably lower CAIR score, not because the metric is less accurate, but because it is difficult to employ and, thereby, not the best contender for a universal privacy metric.
@@
Furthermore, DCR and attribute disclosure have relatively similar scores, but when investigating individual dimensions, it is evident that the two metrics have different downsides. Although DCR performs poorly on Comparability, the attribute disclosure metric has mainly poor scores for Interpretability. This exemplifies that CAIR not only allows stakeholders to rank metrics, but also allows researchers to accurately identify why they perform as they do and where improvements can be made. Interestingly, the same approach can be taken for well-performing metrics. $\epsilon$-identifiability can be improved in C3 data type agnostic by properly accommodating mixed data types and in R1 anomalies by improving the weighting of outliers.
@@
In addition, threshold-based MIA and the attribute disclosure metric share many similarities, but their CAIR scores differ substantially. What is noticeable by investigating Figure 2 is that the Comparability dimensions are scored identically, while the largest difference is in Applicability and a few other dimensions. Based on the difference in Applicability observed in the figure, it can be estimated that attribute disclosure is somewhat more complex compared to threshold-based membership inference attacks.
Overall, these results suggest that the use of a rubric to break down the assessment of privacy evaluation metrics provides detailed insights into strengths and weaknesses of the metrics in comparison and isolation.
6 Discussion and Conclusion
@@
With the increasing awareness of privacy-related issues in data management and sharing, it is more important than ever to focus on streamlining privacy evaluation, especially for maturing research fields such as synthetic data generation. Many domains, especially healthcare, rely heavily on the sharing of data to move the associated research fields forward. Unfortunately, the regulations put in place to protect people decelerate innovations. On the other hand, regulations force institutions to consider the natural right of privacy of people. Therefore, we must work with regulators to best ensure the continuation of potentially lifesaving innovation while still maintaining the privacy of the natural person.
@@
CAIR is an important step in a direction in which regulators and researchers collaborate and enable data sharing. Regulators are starting to notice synthetic data as a possible substitute for true data [6]. Hence, a proper evaluation of the privacy-preserving capabilities of synthetic data is of great importance. A natural first step is to define what constitutes a good universal privacy-evaluation metric for synthetic tabular data. To the best of our knowledge, CAIR is the first attempt to define properties for such measures. Therefore, this work has the potential to guide the research community in a specific direction rather than using arbitrary metrics that researchers find appropriate for their work. Additionally, CAIR serves as guiding principles for regulators in identifying metrics and thresholds for when synthetic tabular data are “private enough” to be shared publicly. In summary, CAIR can help researchers, institutes, and regulators find common ground and agree on what universal privacy metrics are appropriate for synthetic tabular data.
@@
One of the strengths of CAIR is the ability to rank metrics according to the 16 identified dimensions. However, CAIR does not guarantee that a high-scoring metric will be the most accurate metric for any given synthesis or synthetic data. This is not necessarily an issue with the definition of CAIR, but rather with the ideology of having a universal evaluation method. Therefore, we encourage to use universal metrics alongside metrics that are more specific for the given use case. In addition, the dimensions that have been identified as properties for good metrics are likely non-exhaustive. Consequently, future work should look at refining these properties while ensuring alignment with stakeholders. Furthermore, CAIR incentivizes researchers to consider what improvements can be made to existing privacy metrics and to design new metrics that conform to CAIR and the idea of universal evaluation metrics.

### v2

@@ 5.2 Findings
The overall CAIR scores are given by the means with the propagated error and are presented in Figure 5. Of the selected metrics, the most versatile metric for this illustrative example is $\epsilon$-identifiability with a score of $3.44\pm 0.08$. The metric scores well in most dimensions and notably has high scores in all four main categories. In contrast, the most closely related metric to $\epsilon$-identifiability, IDR, significantly worse and is especially penalized in the Applicability principle for requiring information that is difficult to obtain, such as equivalence classes of the true population. Consequently, IDR gets a considerably lower CAIR score, not because the metric is less accurate but because it is difficult to employ and, thereby, not the best contender for a universal privacy metric.
@@
Furthermore, DCR and attribute disclosure receive similar scores for a number of dimensions, but when investigating individual dimensions, it is evident that the two metrics have different downsides. Although DCR performs poorly on Comparability, the attribute disclosure metric mainly has poor scores for Interpretability. This exemplifies that CAIR not only allows stakeholders to rank metrics, but also allows researchers to accurately identify why they perform as they do and where improvements can be made. Interestingly, the same approach can be taken for well-performing metrics. $\epsilon$-identifiability can be improved in C3 data type agnostic by properly accommodating mixed data types and in R1 anomalies by improving the weighting of outliers.
DOMIAS performs the well on Representativeness indicating that the metric generally estimates an accurate privacy level. However, the metric falls behind on Interpretability compared to $\epsilon$-identifiability mainly due to the complex nature of density estimation and the difficulty in disseminating the underlying measurements to non-technical stakeholders. Furthermore, the metric, like all the other metrics, can be improved on C3 data type agnostic by formally facilitating evaluation for data with both discrete and numerical data.
@@
Threshold-based MIA and the attribute disclosure metric share many similarities, but their CAIR scores differ substantially. An examination of Figure 5 reveals that the scores for the Comparability dimension are identical, while the most significant discrepancy lies in the Applicability principle. Based on the difference in Applicability, it can be estimated that attribute disclosure is somewhat more complex compared to threshold-based membership inference attacks.
Interestingly, CAIR score discrepancies can be utilized to identify complementary metrics that, in unison, can provide a comprehensive analysis of the privacy gain for synthetic datasets. Take, for example, DOMIAS and $\epsilon$-identifiability: they share similar scores for both Comparability and Applicability, while the major differences lie in Interpretability and Representativeness. Where DOMIAS excels at accurately representing privacy levels, $\epsilon$-identifiability is easy to interpret. Hence, a comprehensive understanding of privacy can be achieved by employing both methods as complementary evaluation methods. Although this property of CAIR is useful, the primary objective of CAIR remains to identify shortcomings and improve existing or develop new metrics that conform to the CAIR dimensions.
These findings suggest that employing a rubric to systematically scrutinize privacy evaluation metrics yields nuanced and granular insights into their efficacy as universal privacy metrics both in relation to each other and in isolation.
6 Discussion
@@
With the increasing awareness of privacy-related issues in data management and sharing, it is more important than ever to focus on streamlining privacy evaluation, especially for maturing research fields such as synthetic tabular data generation. Many domains rely heavily on the sharing of data to advance the associated research fields. Unfortunately, the regulations put in place to protect people slow down innovation. On the other hand, regulations force institutions to consider the natural right to privacy of people. Therefore, we must work with regulators to best ensure the continuation of potentially lifesaving innovation while still maintaining the privacy of the natural person.
CAIR is an important step in the direction in which regulators and researchers collaborate and enable data sharing. Regulators are beginning to notice synthetic data as a possible substitute for true data (European Comission,, 2021), highlighting the importance of proper evaluation of the privacy preservation capabilities of synthetic data. A natural first step is to define what constitutes a privacy metric that enables accurate privacy estimation while being sufficiently versatile to have regulations established around them. Although precision is the ultimate goal of measuring privacy levels, highly intricate and context-specific measures are difficult to regulate around making attempts at wide implementation futile. Therefore, CAIR serves as a framework for guiding metric development in a direction where precision remains an important aspect while considering outside factors that are crucial for the application of a metric.
@@
To the best of our knowledge, CAIR is the first attempt to define principles for such measures. Therefore, this work has the potential to guide the research community in a specific direction rather than using arbitrary metrics that researchers find appropriate for their work. Furthermore, CAIR serves as guiding principles for regulators in identifying metrics and thresholds that define sufficient privacy levels to publish data.
In summary, CAIR can help researchers, institutes, and regulators find common ground and agree on what universal privacy metrics are appropriate for synthetic tabular data.
6.1 Composition and Application of CAIR
One of the primary objectives of applying CAIR is to identify the strengths and weaknesses of metrics. CAIR enables researchers and other stakeholders alike to analyze how a metric functions and performs, either across the four main principles or the more granular 16 dimensions. While scoring a metric results in a single overall CAIR score, individual dimensions provide arguably more important information that can foster productive discussions and further improvements.
@@ 6.1 Composition and Application of CAIR
Still, the overall CAIR score allows a ranking of metrics according to the 16 identified dimensions. However, the CAIR score primarily reflects the universality of a privacy evaluation metric, rather than its direct quality or precision. Thus, CAIR does not guarantee that a high-scoring metric will be the most accurate metric for any given synthesis or synthetic data. This is not necessarily an issue with the definition of CAIR, but rather with the ideology of having a universal evaluation method. Therefore, we encourage the use of universal metrics alongside metrics that are more specific to the given use case. Complementing a universal metric with one or more context-specific metric(s) allows for a thorough analysis of the privacy landscape within a specific context while being comparable to studies in or outside of the given domain.
Furthermore, the ability to identify complementary privacy evaluation metrics provides added benefits by providing comprehensive coverage of the privacy landscape. However, the primary objective of CAIR is to identify single metrics that can be universally applied to synthetic tabular data generation, potentially with context-specific metrics as a complement. Thus, CAIR should be used to identify strengths and weaknesses of metrics with the intention of improving them rather than combining them. The application of CAIR in this paper suggests that these high-fidelity insights provide valuable information that might otherwise have been overlooked. Thus, CAIR provides a foundation for streamlined development of metrics.
In addition, the dimensions that have been identified as properties for good metrics are likely non-exhaustive. Consequently, future work should look at refining these properties while ensuring alignment with stakeholders. Furthermore, CAIR incentivizes researchers to consider what improvements can be made to existing privacy metrics and to design new metrics that conform to CAIR and the idea of universal evaluation metrics.
6.2 Applicable Domain
The CAIR principles are limited to evaluation of universal privacy metrics for synthetic tabular data and, as such, should not be applied to metrics that fall outside of that domain. Although the CAIR dimensions are designed for tabular data, the four principles (Comparability, Applicability, Interpretability, and Representativeness) may transfer well to other data modalities such as semi-structured or unstructured data, including images and natural language. However, some underlying dimensions likely require modality-specific adaptions to accurately represent challenges in the respective domains. Considering that the Interpretability and Representativeness principles are not concerned with the data format, these principles and their respective dimensions have a high degree of transferability to other data-modalities. Rather than focusing on the effect of the data, the focus is the expected behavior of a good and universal privacy measure.
Although the principles Comparability and Applicability are transferable to other modalities as overarching principles, the underlying dimensions likely require substantial adaptations to modality-specific cases. As an example, a privacy metric for image data should not be concerned with C3 Data type agnostic as images are represented by numerical values. Similarly, while the dimension A3 Performance may be a good fit for other modalities, the scoring definitions in the rubric, Table LABEL:tab:rubric, may require a redefinition to better accommodate domain-specific run-time challenges.
Furthermore, some modalities may require additions or replacements of dimensions to adequately accommodate challenges. As an example on image data, the Representativeness principle could have a dimension that denotes the metric’s ability to align with human visual perception of privacy in images. Accordingly, we allow adding extra dimensions to the principles when defining CAIR for other modalities to the extent that it provides benefits. In such a case, it is important to be aware of the weighting of the overall CAIR score. Either a weighting of each dimension should be defined or each principle should have the same number of dimensions to provide equal weights, as mentioned in Section 3.
6.3 Notions of Privacy
Notions of privacy are changing as technological capabilities are advancing. For example, k-anonymity defines privacy as being indistinguishable from $k-1$ other individuals (Sweeney,, 2002), while differential privacy is concerned with the effect on an algorithmic outcome by adding or subtracting a single individual from a dataset (Dwork and Roth,, 2013). Similarly, privacy capabilities for synthetic data has various notions of privacy. Specifically, membership inference risk (van Breugel et al.,, 2023), attribute disclosure risk (Kaur et al.,, 2021), and similarity-based risk (Guillaudeux et al.,, 2023; Yan et al.,, 2020) are common approaches to defining privacy preserving capabilities of synthetic data. However, these notions are constantly changing as new adversarial capabilities emerge.
CAIR should ideally embrace the dynamic nature of privacy evaluation. Therefore, the four principles are designed to provide a general notion of the universality of privacy evaluation metrics with the underlying dimensions denoting granular properties. Considering that CAIR already encompasses multiple notions of privacy, we argue that it is likely that the four principles will also apply to changing ideas of privacy. On the other hand, it is likely that the underlying dimensions need adaptations or extensions as the research field develops. As discussed earlier, the properties of CAIR are likely non-exhaustive and future refinements are encouraged. This is a consequence of employing a rubric for bias mitigation. The fixed definitions of the rubric values stability over flexibility of the evaluations, making it less adaptable to future development of privacy semantics. However, stability provides better bias mitigation than a vaguely defined rubric.
Lastly, while some methods for introducing privacy, such as k-anonymity (Sweeney,, 2002) and differential privacy (Dwork and Roth,, 2013) allow quantification of the privacy level, they are not evaluation methods and are not limited to synthetic data. Although differential privacy provides formal privacy guarantees and can be used for synthetic data generation, setting the necessary parameters can be challenging (Jordon et al.,, 2021; Yoon et al.,, 2020) and can result in synthetic data that do not conform to other accepted notions of privacy, such as membership inference and attribute disclosure (Sun et al.,, 2023). Consequently, further evaluation is needed to evaluate privacy and allow comparisons between studies. In these cases, metrics with better CAIR scores are a good choice.
6.4 Addressing the Qualitative Nature of CAIR
Due to the qualitative nature inherent in CAIR assessments, there exists the possibility of subjectivity, which can result in discrepancies in scores among different evaluators. However, considering that the primary objective of CAIR is to foster considerations and discussions about privacy metrics, the negative impact of subjectivity diminishes, emphasizing the usefulness and importance of even a single evaluator.
Nevertheless, when comparing CAIR scores across metrics, bias mitigation becomes increasingly important. Although the use of a rubric can mitigate bias to a certain extent, the inherent variability in interpretation calls for additional measures. An effective approach is to incorporate scores from multiple independent evaluators. Aggregating assessments from multiple perspectives minimizes the impact of individual bias and enhances the reliability of the overall assessment process. As the number of evaluators increases, the propagated error decreases, indicating scores that tend to a consensus among evaluators. The error can be interpreted as disagreement among the evaluators, which is an indicator of the level of ambiguity of a metric. In other words, additional evaluators are likely to provide additional insight into the nuances of a metric that makes cross-metric comparisons more objective.
Disagreements in scores can occur in both a single-evaluator setting where a group acts as one evaluator and in a multi-evaluator setting. In the former case, agreements should be resolved through internal discussions that cover the technical aspects of the metric in question. These discussions contribute to a nuanced development of privacy evaluation metrics. Considering disagreements in multi-evaluator settings, discrepancies are more challenging to resolve, especially with large discrepancies. Consequently, the best approach to managing disagreements is to add more independent evaluators and consider the propagated error in the final CAIR score. If one of the evaluators significantly diverges from the others, additional evaluators will lower the error and bring the CAIR score closer to consensus. This is especially important when considering an author of a metric being one of the evaluators.
Furthermore, the incorporation of multiple evaluators incentivizes the research community to collaborate and collectively define good universal metrics while increasing the reliability and trustworthiness of the metrics. This collaborative approach not only enhances the reliability of the assessment but also fosters a more comprehensive understanding of the 16 CAIR dimensions leading to further improvements in CAIR.
Ultimately, researchers, regulators, and natural persons must collectively trust that any published synthetic data maintain the privacy of the individual, which can only be achieved by identifying robust generation methods and evaluating the resulting synthetic data. By evaluating the privacy-preserving capabilities with a universal privacy metric, the development of these generation methods can be accelerated by providing easy comparisons and improving reproducibility of studies.
7 Conclusion
In this work, we have characterized four principles for good universal privacy evaluation metrics for synthetic tabular data consisting of Comparability, Applicability, Interpretability, and Representativeness (CAIR). By dividing each principle into four dimensions, we show that the CAIR rubric offers valuable granular insights into the behavior and utility of metrics, facilitating streamlined evaluation practices across various studies, allowing researchers to effectively improve privacy evaluation methodologies.
Accordingly, CAIR is a tool with multiple applications, including serving as guidelines when designing new metrics, allowing researchers to improve on existing metrics, and providing a tool for regulators to consult when defining how privacy should be evaluated for synthetic tabular data. Although CAIR relies on a qualitative assessment of metrics, it possesses the potential to act as a facilitator for the synthetic data generation community to scrutinize current and new methods to ensure robust evaluation across diverse domains. Thus, the development of the notion of privacy in synthetic data can be refined and act as a catalyst for innovation of generation methods. In addition, we expect that the CAIR principles will foster agreement among researchers and organizations on which universal privacy evaluation metrics are appropriate for synthetic tabular data.
Acknowledgments
This study was funded by Innovation Fund Denmark in the project “PREPARE: Personalized Risk Estimation and Prevention of Cardiovascular Disease”.
Appendix A CAIR Scores with Explanations
The appendix presents explanations of the scores given to the six selected privacy evaluation metrics. Explanations are provided for the 16 dimensions separated into two tables for readability. For succinctness, only comments for Evaluator 1 are presented. Table LABEL:tab:cair_explanation1 presents the first two metrics: $\epsilon$-identifiability, Identity Disclosure Risk (IDR), Table LABEL:tab:cair_explanation2 presents Distance to Closest Record (DCR) and attribute disclosure (Attr. Dis.), and Table LABEL:tab:cair_explanation3 presents threshold-based membership inference attack (T. MIA). and DOMIAS. The CAIR scores for each metric are the mean of dimension means across the evaluators with the propagated error as described in Figure 2.
Table 1: Presentation of CAIR scores for $\epsilon$-identifiability and Identity Disclosure Risk (IDR). The scores by each evaluator are represented as $E_{1}/E_{2}$ for Evaluators 1 and 2, respectively.
Dimensions $\mathbf{E_{1}/E_{2}}$ $\epsilon$-Identifiability $\mathbf{E_{1}/E_{2}}$ IDR
Comparability
C1
Scale
3/4 Mostly invariant, but relies on distance functions without normalization. 3/4 Mostly invariant, but relies on distance functions without normalization.
C2
Metric bounds
4/4 Bounds are $[0,1]$. 4/4 Bounds are $[0,1]$.
C3
Data type agnostic
2/2 Depends on the choice of distance function. Authors suggest Euclidean distance, which performs poorly on categorical data. Consequences can potentially be severe. Remediations can be cumbersome. 2/3 No clear definition of distance function is given. Consequences can potentially be severe. Remediations can be cumbersome.
C4
Cross-domain relevance
4/3.5 Invariant to the domain. 3/2.5 The connection between training data and the population is not necessarily relevant to all domains.
Applicability
A1
Heteroge-neity
3/2 Depends on the choice of distance function(s). The authors suggest Euclidean distance, but explicitly leave the choice of distance function to the user. 3/2 Depends on the choice of distance function(s), but no clear definition is provided.
A2
Diverse generation methods
4/3 Invariant to the generation method. 4/3 Invariant to the generation method.
A3
Performance
3/3 Computes the pairwise distances within the true data and between the true and synthetic data $\mathcal{O}(2n^{2})$. 2/2 Runs in at least $\mathcal{O}(N^{2}+n^{2})$ where $n$ is the sample size and $N$ is population size. $N$ can potentially be extremely large and expensive to estimate.
A4
Implemen-tation
4/3.5 An implementation is provided and the metric is well-documented. 1/1 The metric requires information from the population, which is difficult to obtain or estimate. The error correction parameter $\lambda$ is external information from the literature. The parameter $R$ ”learning something new” is ambiguous and not well-defined and can be challenging to compute.
Interpretability
I1
Explain-ability
4/4 Measures the proportion of synthetic data points that are ”different enough” from true data points. Simple percentage of how many records are at risk. 2/2 Requires knowledge on data errors, sampling, and populations, but can be communicated to someone with some degree of prior knowledge.
I2
Understand-ability
4/3.5 The output is the risk in percentage, which is a common concept for laypeople. 4/3 The output is the risk in percentage, which is a common concept for laypeople.
I3
Visualization
4/3.5 The output is a single value that can easily be visualized. 4/3 The output is a single value that can easily be visualized.
I4
Granularity
4/3 The output is linear with a single value as the output. 3/2 The output is linear with a single value as the output. It is unclear how changes to the output are due to the synthetic data or the population.
Representativeness
R1
Anomalies
3/4 The metric takes all records into account, but the weighting is not definitively defined. The authors suggest using inverse column-wise entropy, but other weightings may be more appropriate. 3/2 The metric is weighted with the inverse size of the equivalence class, but is unclear how appropriate such a weighting is.
R2
Coverage
3/4 Depends on the choice of weightings. The authors suggest inverse column-wise entropy but explicitly leave the choice for the user. 3/2.5 Depending on how the population is computed, the result may not be accurate for all classes.
R3
Reproducibility
4/4 Always produces the same output for the same input. 3/3 Depends on how the population is estimated and the definition of ”learning something new”. Without clear definitions, the results may not be reproducible in all instances.
R4
Precision
3/4 Slightly overestimates privacy violations as the same synthetic data record can pose a violation for an arbitrary number of real records. 3/2.5 Slightly overestimates privacy violations as the same synthetic data record can pose a violation for an arbitrary number of real records.
CAIR $\mathbf{3.44\pm 0.08}$ $\mathbf{2.77\pm 0.09}$
Table 2: Presentation of CAIR scores for Distance to Closest Record (DCR) an attribute disclosure (Attr. Dis.). The scores by each evaluator are represented as $E_{1}/E_{2}$ for Evaluators 1 and 2, respectively.
Dimensions $\mathbf{E_{1}/E_{2}}$ DCR $\mathbf{E_{1}/E_{2}}$ Attr. Dis.
Comparability
C1
Scale
1/1 Completely dependent on the scale of the data. 3/2.5 The metric is largely invariant to the data scale, but normalization is recommended.
C2
Metric bounds
2/2 The metric only has a lower bound $[0,\infty)$. 4/4 Bounds for both outputs are well defined:
Precision: $[0,1]$
Recall: $[0,1]$
C3
Data type agnostic
2/1.5 Depends entirely on the choice of distance function, but without considering how functions are applied, the results can be significantly affected. 2/3 Depends on the choice of distance function, but without considering how functions are applied, the results can be significantly affected.
C4
Cross-domain relevance
2.5/3 Considering some domains use data with a large number of attributes, DCR would not be appropriate for such domains as the distance between records will grow with the number of attributes. For example, in genomics. 3/3 Relevant for most domains as long as the thresholds correspond to the data.
Applicability
A1
Heteroge-neity
3/2.5 Depends on the choice of distance function, which is not for defined for the metric. 3/3 Depends on the choice of distance function, which is not defined for the metric.
A2
Diverse generation methods
4/3.5 Invariant to generation method. 4/4 Invariant to generation method.
A3
Performance
4/3.5 Only uses the pairwise distances. $\mathcal{O}(n^{2})$. 3/2 $\mathcal{O}(kn^{2})$ where k is the number of different levels of knowledge for the adversaries that are tested.
A4
Implemen-tation
3/2 The metric is very simple and easy to implement, but poses some ambiguity. 3/2.5 There is some ambiguity in how predictions are performed and in what order. Additionally, the number of adversarial knowledge levels can vary.
Interpretability
I1
Explain-ability
4/4 Very simple and easy to explain. 2/3 A layperson is unlikely to be familiar with the concepts of prediction and adversarial knowledge. Substantial explanations are required. Some background knowledge is necessary.
I2
Understand-ability
3/4 The concept of distance functions is not necessarily known by laypeople, but can easily be explained. 2/2 The output is precision and recall, which are unfamiliar concepts to laypeople. Without some background knowledge in statistics or data management, it can be difficult to understand.
I3
Visualization
2/3 Comparisons are difficult considering that the metric has no upper bound and the scale is not the same across datasets. 2/2 Visual comparisons become challenging when multiple adversarial levels are being tested unless aggregate data are used (in which case the visualization is no longer completely accurate).
I4
Granularity
1.5/1 Privacy levels can be differentiated, but due to the effect of the scale of the data, some privacy violations may be overlooked. The value only makes sense in comparison to other results. 1/2 The output potentially consists of many sets of values, but at least two values.
Representativeness
R1
Anomalies
2/2 All observations are treated equally. 2/2 There is no specific weighting for data points. A potential weighting should be different for various levels of adversarial knowledge.
R2
Coverage
1/1.5 DCR is only ever representative of small subsets of the data, whether the mean, median, or individual DCRs are considered. 3/3.5 The performance of the predictive attack represents most points, but without considering outliers, some might have overlooked privacy violations.
R3
Reproducibility
4/3.5 Always produces the same output for the same input. 3/2.5 Depends on the implementation and how the prior knowledge is selected.
R4
Precision
1/1.5 When using the median or mean DCR, the privacy level is significantly underestimated. 2/3 Somewhat favors inliers.
CAIR $\mathbf{2.48\pm 0.07}$ $\mathbf{2.69\pm 0.08}$
Table 3: Presentation of CAIR scores for threshold-based MIA (T. MIA) and DOMIAS. The scores by each evaluator are represented as $E_{1}/E_{2}$ for Evaluators 1 and 2, respectively.
Dimensions $\mathbf{E_{1}/E_{2}}$ T. MIA $\mathbf{E_{1}/E_{2}}$ DOMIAS
Comparability
C1
Scale
3/4 The metric is largely invariant to the data scale, but normalization is recommended. 3/3 Mostly invariant, but relies on density estimation methods that may require normalization.
C2
Metric bounds
4/4 Bounds for both outputs are well defined:
Precision: $[0,1]$
Recall: $[0,1]$
4/4 Bounds are [0, 1]
C3
Data type agnostic
2/2 Depends on the choice of distance function, but without considering how functions are applied, the results can be significantly affected. 2/2 Depends on the density estimation method. The authors suggest BNAF or Gaussian KDE, which do not inherently work with categorical data.
C4
Cross-domain relevance
3/3 Relevant for most domains as long as the threshold is appropriate to the domain. 4/4 Invariant to the domain.
Applicability
A1
Heteroge-neity
3/3 Depends on the choice of distance function, which is not defined for the metric. 3/2 Depends on the choice of density estimation og possible underlying distance function.
A2
Diverse generation methods
4/3 Invariant to the generation method. 4/3.5 Invariant to the generation method.
A3
Performance
4/3.5 Computes pairwise distances $\mathcal{O}(n^{2})$. 2.5/2 Depends on the choice of density estimation and possible underlying distance function. BNAF is computationally expensive while Gaussian KDE is $\mathcal{O}(n^{2})$. Accordingly, DOMIAS is at best $\mathcal{O}(2n^{2})$.
A4
Implementation
4/3.5 Easy to compute and implement. 3.5/3.5 An implementation is provided by the authors, but the choice of density estimator is ambiguous.
Interpretability
I1
Explain-ability
3/4 The concept can be explained as “does this synthetic data point look too similar to a true record?”. However, explaining how similarity relates to membership inference can be somewhat difficult to understand for laypeople. 2.5/3 Understanding the use of density estimation likely requires some domain knowledge.
I2
Understand-ability
2/3 The output is precision and recall for an adversarial attack, which are unfamiliar concepts to laypeople. Without some background knowledge in statistics or data management, it can be difficult to understand. 2.5/2.5 Depends on the choice of performance metric. Accuracy is well-understood, while AUROC requires some explanation.
I3
Visualization
4/3 Even though the output consists of two values, they can easily be represented visually. 4/3 The output is a single value that is easy to visualize.
I4
Granularity
1/1.5 The output consists of two values, making differentiation between privacy levels quite challenging. 4/3.5 The output is linear with a single value as the output.
Representativeness
R1
Anomalies
2/2.5 The threshold is the same for all data points. 4/4 Differences in density estimations around outliers are likely greater than for inliers.
R2
Coverage
3/3 The performance of the predictive attack represents most points, but without considering outliers some might have overlooked privacy violations. 3/2.5 The coverage can be affected by the number of features and the corresponding sample size.
R3
Reproducibility
4/3.5 Always produces the same output for the same input. 3/3.5 Depends on the choice of density estimation method.
R4
Precision
1/1.5 Systematically favors inliers by having a fixed similarity threshold. 3/3 Potentially underestimates the privacy level.
CAIR $\mathbf{2.97\pm 0.08}$ $\mathbf{3.16\pm 0.06}$